The perimeter of a rectangle oceanfront lot is 190 meters. The width is one-fourth of the legnth. Find the dimensions.: The perimeter of a rectangle oceanfront lot is 190 meters. The width is one-fourth of the legnth. Find the dimensions.
Answer by HyperBrain(507)

if you remember, the formula for the perimeter is:
P=2(length +width)
The perimeter of a rectangle oceanfront lot is 190 meters.
So, P=190. Thus,
The first equation is
190=2(length+width)
Dividing both sides by 2,
95=length+width
length+width=95----2nd equation

The width is one-fourth of the length.
So, the next equation is:
width=length/4
If you substitute this to the 2nd equation,
length+(length/4)=95
(4*length/4)+(length/4)=95
5*length/4=95
Multiplying both sides by 4,
5*length=380
Dividing both sides by 5,
length=76
Since width=length/4 then,
width=76/4=19

Therefore, the dimensions of the rectangle is 76*19.
Power up,
HyperBrain!